OpenStudy (lilsis76):

what is 90 percent of 80 percent of 70 percent? I feel like I understand it but I really dont. PLEASE HELP!!!

OpenStudy (anonymous):

To start, we need to convert everything into decimals, by diving by 100 So 90/100 = 0.90 80/100 = 0.80 70/100 = 0.70

OpenStudy (lilsis76):

okay, I got that far on one of my scratch papers. @shoplovepc

OpenStudy (anonymous):

So next (well this is what I guess you'd do) is multiply the three so 0.80 * 0.70 * 0.90 = 0.504 0.504 * 100 = 50.4%

OpenStudy (lilsis76):

OH NO WAY!!! I didnt know how to do that haha. Thank you!!! But I have a question. Why do we multiply them all together @shoplovepc

OpenStudy (anonymous):

Well I'm just guessing, honestly. I just sort of figured that when you want to find a percentage of something (ex: 20% of 100) you multiply 100 * 0.20 I just guess I thought that if you want to find the percent of a percent (ex: 50% of 10%) you multiply it by 0.50 * 0.10 and then multiply by 100 :) Tell me if I got it right!

OpenStudy (lilsis76):

YOU DID!! u got it right. I had 5 multiple choice answers and didnt know if my answer was right. thank you MEDAL!!!

OpenStudy (anonymous):

lol yw :)